40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 546     Entry time: Thu Jun 19 20:22:03 2008
Author: rana 
Type: Update 
Category: General 
Subject: FE Computer Status 
I called Rolf (@LLO) who called Alex (@MIT) who suggested that we power cycle every crate
with an RFM connection as we did before (twice in the past year).

Rob and I followed Yoichi around the lab as he turned off and on everything. There
was no special order; he started at the Y-end and worked his way into the corner and
finishing at the X-End. Along the way we also reset the 2 RFM switches around fb0.

This cured the EPICS problem; the FEs could now boot and received the EPICS data.

However, there are still some residual channel hopping-ish issues which Rob and Yoichi are
now working on.
ELOG V3.1.3-